40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 6 17:11:55 2011, steve, Update, SUS, SRM-oplev cable  
    Reply  Wed May 11 11:24:59 2011, steve, Update, SUS, SRM-oplev qpd is back srm-prm_oplevs.jpg
Message ID: 4687     Entry time: Wed May 11 11:24:59 2011     In reply to: 4655
Author: steve 
Type: Update 
Category: SUS 
Subject: SRM-oplev qpd is back 

Quote:

The SRM qpd cable was removed from the BS-table. It's path was changed from 1x4 to ITMY-table following the inner cable tray.

 Laser diode oplev SRM is working. Qpd matrix values were reset like others.

In 0.44mW, returning 0.1mW,  -500 counts.

Attachment 1: srm-prm_oplevs.jpg  135 kB  | Hide | Hide all
srm-prm_oplevs.jpg
ELOG V3.1.3-